OpenStudy (anonymous):

A new movie theater has stadium seating. The first row has 10 seats. Each additional row has 4 seats more than the row before it. How many seats are in the theater if there are 16 rows of seats?

OpenStudy (anonymous):

Formula for partial sums Sn=n/2*(a1+an) a1=10 an=10+4(16-1)=70 n=16 Sn=(16/2)*(10+70) =80*8=640

OpenStudy (campbell_st):

this is an arithmetic sequence 10, 14, 18, ... so to find the number of seats use \[S_{n} = \frac{n}{2}[2a + (n - 1)\times d]\] a is the 1st term in the sequence, n is the number of terms and d is the common difference. In your question you have a = 10, d = 4 and n = 16 substitute and then evaluate, and you'll have the number of seats in the theatre